OpenStudy (anonymous):

All of the items in a bicycle shop are on sale for 30% off the regular price. Rene spends $74.90 to buy a bicycle and a helmet (without taxes). If the regular price of the bicycle is $85, which equation can be used to find the regular price of the helmet, h? a. 0.3(74.90+)=85 b. 0.3(74.90+h)=74.90 c. 0.7(74.90+h)=85 d. 0.7(85+h)=74.90

OpenStudy (anonymous):

All of the items in a bicycle shop are on sale for 30% off the regular price. Rene spends $74.90 to buy a bicycle and a helmet (without taxes). If the regular price of the bicycle is $85, which equation can be used to find the regular price of the helmet, h? a. 0.3(74.90+h)=85 b. 0.3(74.90+h)=74.90 c. 0.7(74.90+h)=85 d. 0.7(85+h)=74.90